Grade 12Algebra

The runs scored in a cricket match by 11 players are as follows: 6, 15, 120, 50, 100, 80, 10, 15, 8, 10, 15 Find the mean, mode, and median of this data. Are the three same?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Arranging the runs scored in a cricket match by 11 players in an ascending order, we get 6, 8, 10, 10, 15, 15, 15, 50, 80, 100, 120
Mean,
Mean of the given data = Sum of all observations/ Total number of observations
= (6 + 8 + 10 + 10 + 15 + 15 + 15 + 50 + 80 + 100 + 120)/ 11
= 429/11
= 39
Mode,
Mode is the value of the variable which occurs most frequently.
Clearly, 15 occurs maximum number of times.
Hence, mode of the given sores is 15
Median,
The value of the middle-most observation is called the median of the data. Here n = 11, which is odd.
Where, n is the number of the students.
∴median = value of 1⁄2 (n + 1)thobservation.
= 1⁄2 (11 + 1)
= 1⁄2 (12)
= 12/2
= 6
Then, value of 6thterm = 15
Hence, the median is 15.
No, these three are not same.

​Thanks